Analyze T-accounts (Learning Objectives 2, 3, 5, & 6) Crystale Displays produces screens for use in various smartphones. The company reports the following information at December 31. Crystale Displays began operations on January 31 earlier that same year. Finished Goods Inventory 127,000 114,000 Work in Process Wages Payable 75,000 75,000 Manufacturing Raw Materials Inventory Inventory Overhead 55,500 36,500 27,500 127,000 64,000 44,000 9,000 44,000 11,000 42,500 Balance 0
Answer #1

1) Direct material used = $27500

2) Indirect material used = 36500-27500 = $9000

3) Cost of direct labor = 64000

4) cost of indirect labor = 75000-64000 = $11000

5) Cost of goods manufactured = $127000

6) Cost of goods sold = 114000

7) Actual manufacturing overhead = 9000+11000+42500 = 62500

8) Manufacturing overhead allocated to jobs = $44000

9) Predetermine overhead rate = 44000/64000 = 68.75% of direct labor

10) Manufacturing overhead is under allocated by (62500-44000) = $18500